# Asymptotic Velocity Domination

### v1

@@ Appendix B Classical Dirac observables
for $\pm{\rm Im}\theta>0$. The occurrance of Bessel functions is explained by the comment after (B.14), where regularity at $\rho=0$ excludes the Bessel $Y$-type solutions. Conversely, it follows from (B) that the Dirac observables (B.18) are a complete set in the sector regular for $\rho/\lambda_{{\textsc{n}}}\rightarrow 0$. Let us stress that (B) holds off-shell, conceptually the Bessel functions entering here are unrelated to those in (2.3).
Defining
${\cal O}_{0}(\theta)$ $:=$ $\Big{(}\frac{\pi}{\lambda_{{\textsc{n}}}\pi_{0}^{\sigma}}\Big{)}^{2}\int_{0}^{2\pi}\!\frac{dx^{1}}{2\pi}\frac{\pi^{\phi}}{\sin^{2}\big{(}\frac{\pi}{\pi_{0}^{\sigma}}(\theta+\tilde{\rho})\big{)}}\,$,
$Q_{1}(\theta)$ $:=$ $\Big{(}\frac{\pi}{\lambda_{{\textsc{n}}}\pi_{0}^{\sigma}}\Big{)}^{2}\int_{0}^{2\pi}\!\frac{dx^{1}}{2\pi}\frac{\rho^{2}\partial_{1}\phi/\lambda_{{\textsc{n}}}^{2}}{\sin^{2}\big{(}\frac{\pi}{\pi_{0}^{\sigma}}(\theta+\tilde{\rho})\big{)}}\,$, (B.21)
one can write (B.18), (B) as follows
${\cal O}(\theta)=\sum_{l\geq 0}\frac{1}{(2l\!+\!1)!}{-3/2\choose l}\Big{(}\frac{-\rho^{2}}{\lambda_{{\textsc{n}}}^{2}}\Big{)}^{l}\Big{(}\frac{\partial}{\partial\theta}\Big{)}^{2l}\Big{[}{\cal O}_{0}(\theta)+\frac{1}{(2l\!+\!2)}\frac{\partial}{\partial\theta}Q_{1}(\theta)\Big{]}\,.$ (B.22)
Recall that $\rho>0$ is a temporal function whose $\rho=0$ level set can be identified with the Big Bang. The leading $\rho=0$ term, with the dynamical variables of the VD system substituted
@@
${\cal O}_{{\textsc{vd}}}(\theta):={\cal O}_{0}(\theta)\big{|}_{\pi^{\phi}\mapsto\pi^{\varphi},\pi^{\sigma}\mapsto\pi^{\varsigma}}\,$, (B.23)
is in fact a one-parameter family of Dirac observables in the VD system. This can be verified directly using the gauge variations (A.10). The subleading terms define an anti-Newtonian expansion around the Carroll limit from Appendix A.2 and come in powers of $\rho^{2}/\lambda_{{\textsc{n}}}^{2}$, inverse powers of the Newton constant in particular. The quantity $Q_{1}(\theta)$ contains the basic spatial derivative needed to make contact to the full Gowdy system. Under the integrals the $\partial/\partial\theta$ derivatives could be replaced with $\partial/\partial\tilde{\rho}$ derivatives, so that (B.22) can alternatively be viewed as providing a spatial gradient expansion of the full Dirac observables around the velocity dominated ones.
@@
On-shell specialization. Subject to the evolution equations the left hand side of the defining relation $\delta_{\epsilon}^{\textsc{h}}\jmath_{0}^{\textsc{p}}(\theta)=\partial_{1}[(\epsilon/n)\jmath_{1}^{\textsc{p}}(\theta)+\epsilon^{1}\jmath_{0}^{\textsc{p}}(\theta)]$ must for $\epsilon=n,\epsilon^{1}=0$ produce the infinitesimal time evolution $e_{0}(\jmath_{0}^{{\textsc{p}}}(\theta))$. The Dirac observables thereby give rise to a one-parameter family of conserved currents
@@
with $T_{n}$ a Wronskian normalized solution of $\partial_{\tau}^{2}+e^{2\tau}n^{2}]T_{n}(\tau)=0$. By construction $\partial_{\tau}{\cal O}^{\rm on}(\theta)=0$. Inserting $\phi^{\rm on}$ into the first line of (B) one finds
@@
In a first step, the coefficients ${\cal O}_{n}(\theta)$ come out as a linear combination of $\partial_{\tau}T_{n}(\tau)$ and $ine^{2\tau}T_{n}(\tau)$ with coefficents that are Fourier coefficients of the $c_{0},c_{1}$. The latter are conveniently evaluated using (B), after which the Wronskian identity for $Y_{1}J_{0}-Y_{0}J_{1}$ yields the above ${\cal O}_{n}(\theta)$ (where $\theta(\mp n)$ is the step function, imposing $\mp n\in\mathbb{N}$). Importantly, for ${\rm Im}\theta\neq 0$ the ${\cal O}_{n}(\theta)$ decay faster than any power in $|n|$. Indeed, for Hadamard states we saw in Section 4.2 that $\lambda_{n}\rightarrow 1$, $\mu_{n}\rightarrow 0$, for $|n|\rightarrow\infty$; the rapid decay would however persist for any $\lambda_{n},\mu_{n}$ polynomially bounded in $|n|$.
An analogous computation can be done for the VD observables (B.23). Here $\varphi^{\rm on}=\sum_{n}\{\mathfrak{t}_{n}(\tau)e^{in\zeta}a_{n}+\mathfrak{t}_{n}(\tau)^{*}e^{-n\zeta}a_{n}^{*}\}$, $\partial_{\tau}^{2}\mathfrak{t}_{n}=0$, needs to be inserted into the much simpler (specialized) on-shell form of ${\cal O}_{0}(\theta)|_{\pi^{\phi}\mapsto\pi^{\varphi},\pi^{\sigma}\mapsto\pi^{\varsigma}}$. The integrals encountered are for $\pm{\rm Im}\theta$ readily evaluated by contour deformation, while $\partial_{\tau}\mathfrak{t}_{n}=(-i/\sqrt{\pi})(\tilde{\lambda}_{n}\!-\!\tilde{\mu}_{n})$ is already $\tau$ independent. The result is
@@
Since by (3.21) $\tilde{\lambda}_{n}\!-\!\tilde{\mu}_{n}=\lambda_{n}\!-\!\mu_{n}$ one arrives at

### v2

@@ Appendix B Classical Dirac observables
for $\pm{\rm Im}\theta>0$. The occurrence of Bessel functions is explained by the comment after (B.14), where regularity at $\rho=0$ excludes the Bessel $Y$-type solutions. Conversely, it follows from (B) that the Dirac observables (B.21) are a complete set in the sector regular for $\rho/\lambda_{{\textsc{n}}}\rightarrow 0$. Let us stress that (B) holds off-shell, conceptually the Bessel functions entering here are unrelated to those in (2.3).
Result B.1.
Set
${\cal O}(\theta):=\int_{0}^{2\pi}\!\frac{dx^{1}}{2\pi}\big\{c_{0}(\tilde{\rho}+\theta,\rho)\pi^{\phi}-c_{1}(\tilde{\rho}+\theta,\rho)\frac{\rho^{2}}{\lambda_{{\textsc{n}}}^{2}}\partial_{1}\phi\big\}\,,\quad{\rm Im}\theta\neq 0\,$, (B.21)
with $c_{0}$, $c_{1}$ given alternatively by (B), (B), (B). Then ${\cal O}(\theta)$ is a one-parameter family of strong, off-shell Dirac observables for the polarized $T^{3}$ Gowdy cosmologies. The integrand $\jmath_{0}^{{\textsc{p}}}(\theta)$ has an ‘anti-Newtonian’ expansion in even powers of $\rho/\lambda_{{\textsc{n}}}$ given in (B.18). The leading term gives rise to one-parameter family of strong, off-shell Dirac observables in the VD system via
@@
${\cal O}^{{\textsc{vd}}}(\theta)=\int_{0}^{2\pi}\!\!\frac{dx^{1}}{2\pi}\,\jmath_{0}^{{\textsc{p}}}(\theta)_{0}\big|_{\pi^{\phi}\mapsto\pi^{\varphi},\pi^{\sigma}\mapsto\pi^{\varsigma}}\,.$ (B.22)
Recall that $\rho>0$ is a temporal function whose $\rho=0$ level set can be identified with the Big Bang. The Dirac observables (B.21) therefore remain regular at the Big Bang and (B.18) provides a systematic expansion away from it, notably in inverse powers of the reduced Newton constant $\lambda_{{\textsc{n}}}$. Upon substitution of the dynamical variables of the VD system the $\rho=0$ term defines Dirac observables in the VD system, as indicated in (B.22). The invariance $\delta_{\epsilon}^{{\textsc{hvd}}}{\cal O}^{{\textsc{vd}}}(\theta)=0$ can be verified by direct computation from (A.10).
@@
On-shell specialization. Subject to the evolution equations the left hand side of the defining relation $\delta_{\epsilon}^{\textsc{h}}\jmath_{0}^{\textsc{p}}(\theta)=\partial_{1}[(\epsilon/n)\jmath_{1}^{\textsc{p}}(\theta)+\epsilon^{1}\jmath_{0}^{\textsc{p}}(\theta)]$ must for $\epsilon=n,\epsilon^{1}=0$ produce the infinitesimal time evolution $e_{0}(\jmath_{0}^{{\textsc{p}}}(\theta))$. The Dirac observables thereby give rise to a one-parameter family of on-shell conserved currents
@@
with $T_{n}$ a Wronskian normalized solution of $\partial_{\tau}^{2}+e^{2\tau}n^{2}]T_{n}(\tau)=0$. By construction $\partial_{\tau}{\cal O}^{\rm on}(\theta)=0$, which can also be verified directly from (B.14). The value of ${\cal O}^{\rm on}(\theta)$ can therefore be found from the $\tau\rightarrow$-$\infty$ limit of the integrand. Since $c_{0}$, $c_{1}$ are regular at $\rho=0$ and AVD implies $\partial_{\tau}\phi\rightarrow v$ for $\tau\rightarrow$-$\infty$ (see Section 2.2) one infers ${\cal O}^{\rm on}(\theta)={\cal O}_{{\textsc{vd}}}^{\rm on}(\theta)$. In fact, a fairly explicit expression for the on-shell value can be found as follows: inserting $\phi^{\rm on}$ into the first line of (B) one finds
@@
We outline the derivation. In a first step, the coefficients ${\cal O}_{n}(\theta)$ come out as a linear combination of $\partial_{\tau}T_{n}(\tau)$ and $ine^{2\tau}T_{n}(\tau)$ with coefficients that are Fourier coefficients of the $c_{0},c_{1}$. The latter are conveniently evaluated using (B) and come out as
$\int_{0}^{2\pi}\!\frac{d\zeta}{2\pi}e^{in\zeta}c_{0}(\zeta+\theta,e^{\tau})=-J_{0}(e^{\tau}|n|)e^{\pm i|n|\theta}\theta(-|n|)\,$,
$\int_{0}^{2\pi}\!\frac{d\zeta}{2\pi}e^{in\zeta}c_{1}(\zeta+\theta,e^{\tau})=\pm i|n|J_{1}(e^{\tau}|n|)e^{\pm i|n|\theta}\theta(-|n|)\,$, (B.26)
where $\theta(\mp n)$ is the step function, imposing $\mp n\in\mathbb{N}$. Expressing $\partial_{\tau}T_{n}(\tau)$ and $ine^{2\tau}T_{n}(\tau)$ in terms of Bessel $J_{1},Y_{1}$ and $J_{0},Y_{0}$ functions via (2.3) many terms cancel and use of the Wronskian identity $Y_{1}(x)J_{0}(x)-Y_{0}(x)J_{1}(x)=-2/(\pi x)$ yields the above ${\cal O}_{n}(\theta)$. Importantly, for ${\rm Im}\theta\neq 0$ the ${\cal O}_{n}(\theta)$ decay faster than any power in $|n|$. Indeed, for Hadamard states we saw in Section 4.2 that $\lambda_{n}\rightarrow 1$, $\mu_{n}\rightarrow 0$, for $|n|\rightarrow\infty$; the rapid decay would however persist for any $\lambda_{n},\mu_{n}$ polynomially bounded in $|n|$.
An analogous computation can be done for the VD observables (B.22). Here $\varphi^{\rm on}=\sum_{n}\{\mathfrak{t}_{n}(\tau)e^{in\zeta}a_{n}+\mathfrak{t}_{n}(\tau)^{*}e^{-n\zeta}a_{n}^{*}\}$, $\partial_{\tau}^{2}\mathfrak{t}_{n}=0$, needs to be inserted into the much simpler (specialized) on-shell form of ${\cal O}_{0}(\theta)|_{\pi^{\phi}\mapsto\pi^{\varphi},\pi^{\sigma}\mapsto\pi^{\varsigma}}$. The integrals encountered are for $\pm{\rm Im}\theta>0$ readily evaluated by contour deformation, while $\partial_{\tau}\mathfrak{t}_{n}=(-i/\sqrt{\pi})(\tilde{\lambda}_{n}\!-\!\tilde{\mu}_{n})$ is already $\tau$ independent. The result is
@@
Since by (3.21) $\tilde{\lambda}_{n}\!-\!\tilde{\mu}_{n}=\lambda_{n}\!-\!\mu_{n}$ one confirms